Could someone help me with this word problem please ?
Solve by any method and show equations and major steps. A student has a total of 225 points on three tests. The sum of the scores on the first and second tests exceeds the score on the third test by 61. The first score exceeds the second by 6. Find the three scores

Found 2 solutions by mukhopadhyay, Paul:
Answer by mukhopadhyay(490)

Let x be the first score.
The score on the second test is (x-6)
The score on the third test = x+(x-6)-61
Total score = 225
x+(x-6)+[x+(x-6)-61] = 225
=> 2x+2x-12-61 = 225
=> 4x = 298
=> x = 74.5
Scores on the test are as below:
First = 74.5, Second = 68.5, Third = 82

Answer by Paul(988)

First score is x+6
Second score is x
Third score is y
Equation 1:
x+6+x+61=y
y=2x+67 (subsitution)
Equation 2:
x+x+6+y=225
2x+6+2x+67=225
4x=152
x=38
38+6=44
38(2)+67=143
Hence, the score on the first test was 44, on the second test was 38 and on the third test was 143.
Paul.